OpenStudy (anonymous):

radical 8 + radical 50

OpenStudy (hunus):

\[\sqrt{8} + \sqrt{50} = \sqrt{2^3} + \sqrt{25*2} = \sqrt{2^3} + \sqrt{5^2*2}\] Does that help at all?